OpenStudy (anonymous):

Please help asap, will medal! Simplify the rational expressions. state and excluded values

OpenStudy (anonymous):

\[\frac{ 2x-8 }{ x-4 }\]

OpenStudy (cwrw238):

look at the top part ans see if you can factor it what number divides into 2x and 8 ?

OpenStudy (anonymous):

2 or no because of the x

OpenStudy (cwrw238):

2 is correct

OpenStudy (cwrw238):

so 2x - 8 = 2(x - ?) ? = what

OpenStudy (anonymous):

4?

OpenStudy (cwrw238):

8 divided by 2 = ?

OpenStudy (cwrw238):

right so the top part is 2( x - 4) so we have